OpenStudy (nerdgamer_):

Which value is the greatest? the fraction three-fifths the fraction seven-eighths the decimal seventy-five hundredths the decimal eight tenths - I think its this or b

OpenStudy (nerdgamer_):

I will give medal to whoever helps me

OpenStudy (nerdgamer_):

and a fan

OpenStudy (anonymous):

The fraction three-fifths = 3/5 = .60 the fraction seven-eights = 7/8 = .875 the decimal seventy-five hundredths = .75 the decimal eight tenths =.8

OpenStudy (anonymous):

Which one looks bigger?

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

There's your answer..

OpenStudy (nerdgamer_):

ok thxs soo much

OpenStudy (anonymous):

Haha, no problem :)